OpenStudy (anonymous):

simplify log2 (8)^3.5

OpenStudy (anonymous):

log2(2)^3*3.5 log2(2)^10.5 10.5log2(2)=10.5*1=10.5

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

simplify: log9 (1/81)^5